# Enhancing Architecture Frameworks by Including Modern Stakeholders and their Views/Viewpoints

### v3

@@ I Introduction
The contribution of this paper is the identification of ML, data science, and data engineering stakeholders and their concerns, as well as the division of modeling perspectives, viewpoints, and views that can frame and address those concerns to enable a sophisticated and thorough architecture description for ML-enabled systems.
@@ V-B Proposed viewpoints and views
According to our own knowledge about the state of practice in the field of DEA (or DSE), we envisioned the mathematical notations commonly used in DA and ML, as well as the graphical notation of PGMs [26], and the Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) [18] for ANNs, as potentially adequate candidates for providing the necessary model kinds that should serve the architecture views required for supporting the proposed new architecture viewpoints. In addition, we assumed that existing UML diagrams could be adopted for adapting the viewpoints of other stakeholders in order to enable their appropriate communication and collaboration with data scientists and data engineers. The empirical study through the survey questionnaire confirmed these assumptions and helped us devise new viewpoints and views using existing notations and model kinds, as illustrated in Table V. The list of notations and model kinds in each row of the table is ordered based on the opinions of the survey participants concerning the suitability of each option for the specific purpose. The radar diagrams in Figures 2 and 3 show this for the collaboration of data scientists with their peers, whereas the ones in Figures 4 and 5 demonstrate this for the collaboration of data engineers with their peers. In fact, we asked the same question from each stakeholder group in two different ways: what is the best model kind and notation for their collaboration; and what is the most suitable model kind and notation for describing the system architecture from their viewpoint?
@@
Figure 2: Model kinds and notations for the collaboration of data scientists with their peers. Choices (clockwise): a) Computational Graphs (CG), also known as Data-Flow Graphs (DFG). b) UML class diagrams. c) Mathematical notation showing the mathematical model (e.g., probability distributions, as well as objective or loss function). d) Data flow diagrams (e.g., using the UML Activity diagram notation) showing the upstream and downstream components in the pipeline. e) Charts, diagrams, or plots (e.g., histograms, pie charts, and scatter plots). f) Entity-Relationship (ER) diagrams. g) Probabilistic Graphical Models (PGMs). h) Topic maps, knowledge graphs (e.g., RDF graphs), or Ontologies. i) Others (please specify).
Figure 3: Architecture description from the perspective of data scientists. Choices (clockwise): i) The computational operations and the flow of data among them. ii) The processes or components and the flow of data among them. iii) The processes or components and the flow of control among them. iv) The underlying mathematical model. v) The data visualization. vi) The workflow or pipeline for data analytics and machine learning. vii) The problem (use case) domain concepts and their relationships. viii) Others (please specify).
Figure 4: Model kinds and notations for the collaboration of data engineers with their peers. The choices were the same as in Figure 2.
Figure 5: Architecture description from the perspective of data engineers. The choices were the same as in Figure 3.
@@ VII Conclusion and future work
Data availability
To promote the open science policy and enable transparency and replicability, we offer the supplementary material of this paper as open data on Zenodo: https://doi.org/10.5281/zenodo.12532455.
@@ Acknowledgments
The authors would like to thank the following interview and survey participants, as well as the anonymous ones, for dedicating their time to this study. Further, they are grateful for the valuable feedback of the anonymous reviewers and the editors, which improved the quality of this paper. Here is the list of names with no specific order: Sebastian Eder (Qualicen GmbH, Germany), Pragya Kirti Gupta (fortiss GmbH, Germany), Bertrand Charpentier (Technical University of Munich, Germany), Nicholas Gao (Technical University of Munich, Germany), Günter Neumann (DFKI GmbH, German Research Center for AI, Germany), Shreya Desai (University of Reading, UK), Alper Kanak (ERARGE Co. Ltd., Turkey), Sander Vanhove, Cláudio Gomes (Aarhus University, Denmark), Hussein Marah (University of Antwerp, Belgium), Raheleh Eslampanah (University of Antwerp, Belgium), Onur Kilincceker (University of Paderborn, Germany and Mugla Sitki Kocman University, Turkey), Dušan Savić (Belgrade University, Serbia), and Ivan Ruchkin (University of Pennsylvania, USA).

### v4

@@ I Introduction
The contribution of this paper is twofold: (i) Identifying ML, data science, and data engineering stakeholders and their concerns; (ii) Proposing modeling perspectives, viewpoints, and views that can frame and address those concerns to enable a sophisticated and thorough architecture description for ML-enabled systems.
@@ V-B Proposed viewpoints and views
According to our own knowledge about the state of practice in the field of DEA (or DSE), we envisioned the mathematical notations commonly used in DA and ML, as well as the graphical notation of PGMs [26], and the Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) [18] for ANNs, as potentially adequate candidates for providing the necessary model kinds that should serve the architecture views required for supporting the proposed new architecture viewpoints. In addition, we assumed that existing UML diagrams could be adopted for adapting the viewpoints of other stakeholders in order to enable their appropriate communication and collaboration with data scientists and data engineers. The empirical study through the survey questionnaire confirmed these assumptions and helped us devise new viewpoints and views using existing notations and model kinds, as illustrated in Table V. The list of notations and model kinds in each row of the table is ordered based on the opinions of the survey participants concerning the suitability of each option for the specific purpose. Tables VI and VII show this from the viewpoint of data scientists, whereas Tables VIII and IX illustrate this from the viewpoint of data engineers. In fact, we asked the same question from each stakeholder group in two different ways: (i) What is the best model kind and notation for their collaboration? (ii) What is the most suitable model kind and notation for describing the system architecture from their viewpoint?
@@
TABLE VI: Notations and model kinds for the collaboration of data scientists with their peers
Notation/model kind Rank based on the votes
Mathematical notation showing the mathematical model (e.g., probability distributions, as well as objective or loss function) 1st
Charts, diagrams, or plots (e.g., histograms, pie charts, and scatter plots) 2nd
Computational Graphs (CG), also known as Data-Flow Graphs (DFG) 3rd
Probabilistic Graphical Models (PGMs) 4th
Data flow diagrams (e.g., using the UML Activity diagram notation) showing the upstream and downstream components in the pipeline 5th
Topic maps, knowledge graphs (e.g., RDF graphs), or Ontologies 6th
UML class diagrams 7th
Entity-Relationship (ER) diagrams 8th
TABLE VII: Notations and model kinds for the software system architecture description from the viewpoint of data scientists
Notation/model kind Rank based on the votes
The workflow/pipeline for data analytics and machine learning 1st
The underlying mathematical model 2nd
The computational operations and the flow of data among them 3rd
The problem/use case domain concepts and their relationships 4th
The data visualization (e.g., in the case of online learning/stream processing) 5th
The processes and/or components and the flow of data among them 6th
The processes and/or components and the flow of control among them 7th
TABLE VIII: Notations and model kinds for the collaboration of data engineers with their peers
Notation/model kind Rank based on the votes
Data flow diagrams (e.g., using the UML Activity diagram notation) showing the upstream and downstream components in the pipeline 1st
UML class diagrams 2nd
Computational Graphs (CG), also known as Data-Flow Graphs (DFG) 3rd
Mathematical notation showing the mathematical model (e.g., probability distributions, as well as objective or loss function) 4th
Entity-Relationship (ER) diagrams 5th
Charts, diagrams, or plots (e.g., histograms, pie charts, and scatter plots) 6th
Topic maps, knowledge graphs (e.g., RDF graphs), or Ontologies 7th
Probabilistic Graphical Models (PGMs) 8th
TABLE IX: Notations and model kinds for the software system architecture description from the viewpoint of data engineers
Notation/model kind Rank based on the votes
The processes and/or components and the flow of data among them 1st
The computational operations and the flow of data among them 2nd
The processes and/or components and the flow of control among them 3rd
The workflow/pipeline for data analytics and machine learning 4th
The problem/use case domain concepts and their relationships 5th
The data visualization (e.g., in the case of online learning/stream processing) 6th
The underlying mathematical model 7th
@@ Acknowledgments
The authors would like to thank the following interview and survey participants, as well as the anonymous ones, for dedicating their time to this study. Further, they are grateful for the valuable feedback of the anonymous reviewers and the editors, which improved the quality of this paper. Here is the list of names of the interview partners and survey participants, who opted to share their names, with no specific order: Sebastian Eder (Qualicen GmbH, Germany), Pragya Kirti Gupta (fortiss GmbH, Germany), Bertrand Charpentier (Technical University of Munich, Germany), Nicholas Gao (Technical University of Munich, Germany), Günter Neumann (DFKI GmbH, German Research Center for AI, Germany), Shreya Desai (University of Reading, UK), Alper Kanak (ERARGE Co. Ltd., Turkey), Sander Vanhove, Cláudio Gomes (Aarhus University, Denmark), Hussein Marah (University of Antwerp, Belgium), Raheleh Eslampanah (University of Antwerp, Belgium), Onur Kilincceker (University of Paderborn, Germany and Mugla Sitki Kocman University, Turkey), Dušan Savić (Belgrade University, Serbia), and Ivan Ruchkin (University of Pennsylvania, USA).